40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 1 20:20:15 2016, gautam, Update, IMC, IMC loss measurement plan IMG_3471.JPG
    Reply  Fri Dec 2 16:40:29 2016, gautam, Update, IMC, 24V fuse pulled out 
       Reply  Mon Dec 5 01:58:16 2016, gautam, Update, IMC, IMC ringdowns IMCringdown.pdfIMCringdown_2.pdf
          Reply  Mon Dec 5 15:55:25 2016, gautam, Update, IMC, IMC ringdowns 7x
             Reply  Mon Dec 5 19:29:52 2016, gautam, Update, IMC, IMC ringdowns REFLanomaly.pdf
                Reply  Tue Dec 6 00:43:41 2016, gautam, Update, IMC, more IMC ringdowns 
                   Reply  Wed Dec 7 11:52:48 2016, ericq, Update, IMC, Partial IMC ringdowns IMCpartial.pngIMCpartial.zip
                      Reply  Thu Dec 8 19:01:21 2016, rana, Update, IMC, Partial IMC ringdowns 
                   Reply  Sun Dec 2 17:26:58 2018, gautam, Update, IMC, IMC ringdown fitting tauTheoretical.pdfringdownFit.pdfringdownScatter.pdfcavPole.pdf
                      Reply  Fri Dec 7 12:51:06 2018, gautam, Update, IMC, IMC ringdown fitting time_reflAndTrans.pdfcorner_transOnly.pdfcorner_reflAndTrans.pdf
Message ID: 12663     Entry time: Mon Dec 5 01:58:16 2016     In reply to: 12660     Reply to this: 12665
Author: gautam 
Type: Update 
Category: IMC 
Subject: IMC ringdowns 

Over the weekend, I worked a bit on getting these ringdowns going. I will post a more detailed elog tomorrow but here is a quick summary of the changes I made hardware-wise in case anyone sees something unfamiliar in the lab...

  • PDA10CF PD installed on PSL table in the beam path that was previously used for the 3f cancellation trials
  • PDA255 installed on MC2 trans table, long BNC cable running from there to vertex via overhead cable tray
  • PDA255 installed on AS table in front of one of the (currently unused) WFS

I spent a while in preparation for these trials (details tomorrow) like optimizing AOM alignment/diffracted power ratio, checking AOM and PMC switching times etc, but once the hardware is laid out, it is easy to do a bunch of ringdowns in quick succession with an ethernet scope. Tonight I did about 12 ringdowns - but stupidly, for the first 10, I was only saving 1 channel from the oscilloscope instead of the 3 we want to apply the MIT method.

Here is a representative plot of the ringdown - at the moment, I don't have an explanation for the funky oscillations in the reflected PD signal, need to think on this.. More details + analysis to follow...


Dec 5 2016, 130pm:

Actually the plot I meant to put up is this one, which has the time window acquired slightly longer. The feature I am referring to is the 100kHz oscillation in the REFL signal. Any ideas as to what could be causing this?

Attachment 1: IMCringdown.pdf  49 kB  | Hide | Hide all
IMCringdown.pdf
Attachment 2: IMCringdown_2.pdf  57 kB  Uploaded Mon Dec 5 13:40:51 2016  | Hide | Hide all
IMCringdown_2.pdf
ELOG V3.1.3-